OpenStudy (anonymous):

What is the quotient in simplest form? State any restrictions on the variable. x+2/ x-1 divided by x + 4/x^2 + 4x - 5

OpenStudy (anonymous):

\[\frac{ x+2}{ x-1 }\div \frac{ x+4 }{ x^2+4x-5 }=\frac{ x+2 }{ x-1 }\div \frac{ x+4 }{ (x+5)(x-1) }\] \[=\frac{ x+2 }{ x-1 }\times \frac{(x+5)(x-1) }{x+4 }\] Simplify it now.

OpenStudy (anonymous):

(x+2)(x+5)/x+4 ?

OpenStudy (anonymous):

what are the restrictions?

OpenStudy (anonymous):

you there @Azteck

OpenStudy (anonymous):

What can't x equal to?

OpenStudy (anonymous):

What number do you prevent x to be so that the denominator of the fraction does not become undefined?

OpenStudy (anonymous):

\[\dfrac{(x+2)(x+5)}{x+4}\] What can't x be???

OpenStudy (anonymous):

@algebra2sucks

OpenStudy (anonymous):

1,-4,-5?

OpenStudy (anonymous):

Dude, if x=1, will there be a value coming out from that fraction there?

OpenStudy (anonymous):

I'm not so sure, but I reckon they're asking you for restrictions after you simplify the quotient.

OpenStudy (anonymous):

Because questions don't ask for you to do something in the beginning at the end of the question. They rather put the question and command you to do stuff in chronological order...

OpenStudy (anonymous):

Ooooh ok so its just -5, and -4??

OpenStudy (anonymous):

Is that right @Azteck

OpenStudy (anonymous):

Why -5?

OpenStudy (anonymous):

Put -5 into that fraction. What do you get? Do you undefined? If you don't get undefined, then -5 ain't restricting noone.

OpenStudy (anonymous):

Well these are the answers I have A. (x+2)(x+5)/x+4; x = -5, -4 B. (x+2)(x+4)/(x+4)^2(x+5); x = 1, -5 C. (x+2)(x+4)/(x+4)^2(x+5); x = 1, -5,-4 D. (x+2)(x+5)/x+4; x = 1, -4, -5

OpenStudy (anonymous):

Then you have to find the restrictions in the quotient. So you were correct before. D is the answer. These questions are dodgy, if they didn't give you multiple choice. Next time write the WHOLE QUESTION down before you post...